# An Evolutionary Algorithm to Accelerate High-Dimensional Optimization

### v2

@@ Abstract.
High-dimensional numerical optimization presents a persistent challenge. This paper introduces Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), an evolutionary algorithm to accelerate convergence in complex, non-differentiable problems afflicted by the curse of dimensionality.
Evaluated on the notoriously difficult CEC2017 benchmark suite of 29 functions, QUASAR achieved the lowest overall rank sum (150) using the Friedman test, significantly outperforming L-SHADE (229) and standard DE (305) in the dimension-variant trials. QUASAR also proves computationally efficient, with run times averaging $1.4\times$ faster than DE and $7.8\times$ faster than L-SHADE ($p\ll 0.001$) in the population-variant trials.
Building upon Differential Evolution (DE), QUASAR introduces a highly stochastic architecture to dynamically balance exploration and exploitation. Inspired by the probabilistic behavior of quantum particles in a stellar core, the algorithm implements three primary components that augment standard DE mechanisms: 1) probabilistically selected mutation strategies and scaling factors; 2) rank-based crossover rates; 3) asymptotically decaying reinitialization that leverages a covariance matrix of the best solutions to introduce high-quality genetic diversity.
QUASAR’s performance establishes it as an effective, user-friendly optimizer for complex high-dimensional problems.
@@ 1. Introduction
High-dimensional numerical optimization remains a persistent challenge across diverse fields, including machine learning, science, engineering, and finance (zhou2024evolutionary). This challenge stems from the curse of dimensionality; the exponential growth of the search space renders gradient-based and parametric search methods ineffective (shan2008survey).
Consequently, optimization metaheuristics, particularly evolutionary algorithms, have become essential for locating optimal solutions in complex function landscapes.
Among these methods, Differential Evolution (DE) has established itself as a robust and highly effective optimizer due to its simplicity, efficiency, and effective mutation strategies (storn1996usage). However, standard DE variants can struggle to maintain an effective balance between global exploration and local exploitation, particularly in high-dimensional search spaces. This often leads to stagnation or premature convergence, requiring frequent restarts or extensive hyperparameter tuning; a significant practical drawback for real-world applications.
To address these limitations, this paper proposes Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), an evolutionary algorithm designed to dynamically balance the exploration-exploitation trade-off and achieve superior performance in higher dimensions. This method is conceptually inspired by the search for stability in a stellar core, where quantum particles probabilistically explore configurations searching for the most stable state.
@@
The efficacy of QUASAR is demonstrated through rigorous testing against standard DE and its state-of-the-art variant L-SHADE across the complete CEC2017 benchmark suite. The results show that QUASAR achieves large and statistically significant performance gains, achieving an overall Friedman rank sum of 150, far below L-SHADE (229) or DE (305). These improvements are achieved using default parameters, highlighting QUASAR’s ease of use.
The remainder of this paper is organized as follows: Section 2 details the QUASAR algorithm, including its mutation, crossover, selection, and reinitialization mechanisms. Section 3 defines the experimental design, and Section 4 discusses the experimental results compared to DE and L-SHADE. Section 5 details the ease of use and implementation of the algorithm. Section 6 concludes the work and Section 7 suggests future work.
@@ 2. QUASAR Algorithm
Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR) operates using a highly stochastic, multi-strategy approach.
@@ 2.1. Initial Population
The initial population is generated as a Sobol sequence by default, due to its low discrepancy and superior uniformity in high-dimensional spaces compared to other Quasi-Monte Carlo sampling methods (bonneel2025sobol). The default population size, $N$, is set to $10\times D$ (where $D$ is the number of dimensions).
Additional initialization methods are available, including uniform Latin Hypercube and Random distributions, as well as exploitative Hyperellipsoid Density sequences (hyperellipsoid).
@@ 2.2. Mutation
Solutions are recombined with one another, analogous to particles sharing information via quantum entanglement. This ’spooky action at a distance’ (A. Einstein, 1947), serves as the conceptual basis for the mutation mechanisms below.
QUASAR uses three distinct mutation strategies, each being a modified Differential Evolution variant. These strategies are selected dynamically based on the specified entanglement probability, entanglerateentangle\ rate. This value defaults to $0.33$, resulting in the three strategies being applied equally across the population.
2.2.1. Local Mutation (Exploitation)
The local mutation strategy, ’Spooky-Best’, is chosen with probability entanglerateentangle\ rate. This strategy is similar to the $\text{DE}/\text{best}/1$ format, exploiting the immediate search space around the current best solution $\mathbf{X}_{\text{best},g}$.
@@ 2.2.1. Local Mutation (Exploitation)
2.2.2. Global Mutations (Exploration)
Two global mutation strategies are applied to the remaining (1entanglerate)(1-entangle\ rate) fraction of the population. Promoting global exploration, these strategies, ’Spooky-Current’ and ’Spooky-Random’, each have a 50% probability of being selected:
@@ 2.2.2. Global Mutations (Exploration)
  • $\mathbf{v}_{i,g}$: The mutant vector for the $i$-th individual at generation $g$.
  • $\mathbf{X}_{i,g}$: The current solution vector.
  • $\mathbf{X}_{\text{best},g}$: The best solution vector known in the population.
  • $\mathbf{X}_{\text{rand},g}$: A randomly chosen solution vector.
  • $F_{\text{local}}$ and $F_{\text{global}}$: The corresponding mutation factors.
@@ 2.3. Crossover
QUASAR uses binomial crossover with a dynamic crossover rate ($CR_{i}$), inversely proportional to the solution’s fitness rank. This mechanism ensures that unstable solutions have a higher chance of inheriting components from the new mutant vector $\mathbf{v}_{i}$.
The raw crossover rate ($CR_{raw},i$) for the $i$-th solution is calculated based on its rank ($rank_{i}$), where $rank=1$ corresponds to the best solution (lowest fitness). The final dynamic crossover rate $CR_{i}$ is capped at a minimum of $0.33$:
(1) $CR_{raw,i}=\frac{\text{max\_rank}-\text{rank}_{i}}{\text{max\_rank}}=\frac{N-1-\text{rank}_{i}}{N-1}$
(2) $CR_{i}=\max(CR_{raw,i},0.33)$
@@ 2.4. Selection
A greedy elitism strategy is used to determine the next generation’s population, where the target vector for generation $g+1$ is the better of the two solutions between the current vector $\mathbf{X}_{i,g}$ and the trial vector $\mathbf{u}_{i,g}$:
@@ 2.5. Asymptotic Reinitialization
The covariance-guided vector reinitialization applies to a fixed 33% of the worst-performing solutions. These solutions have a high chance in early generations of ’tunneling’ to more promising regions.
@@
(5) $P_{\text{reinit}}(g)=\exp\left(\frac{\ln(P_{{}_{\text{final}}})}{g_{{}_{\text{final}}}g_{{}_{\text{max}}}}\cdot g\right)$
where $g_{\max}$ is the maximum generation count, $P_{\text{final}}=0.33$ is the target final probability, and $g_{final}=0.33$ defines the fraction of generations at which $P_{reinit}$ reaches $P_{final}$. This ensures the probability drops to $P_{\text{final}}$ at $g=g_{\max}\cdot g_{\text{final}}$.
When a reinitialization occurs, new positions are sampled, informed by the covariance of top-performing individuals. The best $M$ solutions (where $M=25\%\text{ of the population}$) are used to calculate the mean vector $\boldsymbol{\mu}$ and the covariance matrix $\boldsymbol{\Sigma}$:
@@
(7) $\boldsymbol{\Sigma}=\frac{1}{M-1}\sum_{i=1}^{M}(\mathbf{x}_{i}-\boldsymbol{\mu})(\mathbf{x}_{i}-\boldsymbol{\mu})^{\text{T}}+\epsilon\mathbf{I}$
where $\epsilon$ is a small scalar ($10^{-12}$) and $\mathbf{I}$ is the identity matrix, added to prevent $\boldsymbol{\Sigma}$ from being singular.
@@
The noise vector is $\boldsymbol{\delta}$ is sampled from a multivariate normal distribution, with respect to the search bounds:
(9) $\boldsymbol{\delta}\sim\mathcal{N}\left(\mathbf{0},\left(\frac{\mathbf{b_{high,n}}-\mathbf{b_{low,n}}}{20}\right)^{2}\mathbf{I}\right)$
@@
(10) $\mathbf{X_{new,n}}=\text{clip}(\mathbf{y_{n}},\mathbf{B_{n}})$
@@
Due to the $\mathcal{O}(N)$ scaling of the covariance matrix calculations, along with decaying probability of occurrence, the additional overhead becomes negligible compared to the objective function evaluations.
@@ 3. Experimental Design
3.1. Comparison Optimizers & Test Functions
@@
This choice is justified by the fact that the QUASAR algorithm is built on the DE framework, as well as DE’s standing as a robust and effective optimizer (price1996differential). L-SHADE was specifically chosen due to its success in CEC optimization competitions (mo2019efficient).
@@ 3.2. Experimental Setup
To ensure a fair comparison, each optimizer was run using their default hyperparameters. The initial population generation methods were also left to defaults; the tested sample sizes not being powers of 2 diminishes the uniformity of QUASAR’s initial Sobol sequence (joe2008constructing), providing a theoretical initial handicap.
Since SciPy’s DE implementation internally scales the input population size by dimensionality ($popsize\times D$ when all dimensions have equal bounds), a correction factor of $1/D$ is applied to its initial sample size.
The computational budget for metaheuristic comparison is often defined by the maximum number of objective function evaluations ($FES_{max}$) (kazikova2021objective). Due to L-SHADE’s linear population reduction, the $FES_{max}$ becomes a difficult comparison metric to compare against. As such, the primary comparisons are performed on final solution fitness errors and overall optimization times, given the same number of generations, using modern implementations via SciPy’s DE and MealPy’s L-SHADE.
By default, the L-SHADE implementation features tolerance parameters for early convergence, and the QUASAR and DE implementations further include a final polishing step using L-BFGS-B minimization. These were all disabled for the trials to ensure the optimizers fully utilize their computational budget, terminating only when reaching the maximum number of generations $g_{max}$.
@@ 3.3. Experimental Trials:
The results are compared based on the overall aggregated rank sums, geometric mean error factors ($GMERF$), and optimization run times.
Dimension-Variant Trials:
These trials assessed QUASAR’s scalability by varying the problem dimension $D$ with a fixed population size ($N=1000$).
The dimension-variant experiments were run with $30$ independent trials for each function/dimension combination to ensure statistical significance.
The dimensions were varied between $[10,30,50,100]$, resulting in a total of 10,400 full optimization runs; 3,480 for each algorithm. The computational budget was kept constant, with maximum number of iterations $g_{max}=100$ and sample size $N=1000$.
Table 1. Dimension-Variant Test Parameters
Parameter Test Values
Dimensions ($D$) [10, 30, 50, 100]
Population Size ($N$) 1000
Generations ($g_{\max}$) 100
Sample-Variant Trials:
These trials assessed the impact of population size on performance in high dimensions (a fixed $D=100$). The computational budget is varied between population sizes of [100, 250, 500, 1000, 5000], with 30 independent trials performed for each sample size / test function combination using maximum number of iterations $g_{max}=100$.
The $N=5000$ trials, however, were only run for 10 trials to ensure statistical significance while mitigating the high computational overhead, resulting in a total of 11,310 full optimization runs; 3770 for each algorithm.
Table 2. Sample-Variant Test Parameters
Parameter Test Values
Dimensions ($D$) 100
Population Size ($N$) [100, 250, 500, 1000, 5000]
Generations ($g_{\max}$) 100
@@ 3.4. Statistical Analysis:
For each unique combination, the median final error value ($f$) across all independent trials was calculated for each optimizer. Optimizers were ranked based on these median errors, with the best performance (lowest error) receiving a rank of 1.
@@
Solution Quality Improvements:
The primary comparison metric used to quantify the magnitude of performance difference is the Geometric Mean Error Reduction Factor $GMERF$, which provides an unbiased measure of relative performance across all $N_{\text{trials}}$ independent runs for a given scenario:
$\text{GMERF}=\left(\prod_{i=1}^{N_{\text{trials}}}\frac{f_{\text{comparison},i}}{f_{\text{QUASAR},i}}\right)^{\frac{1}{N_{\text{trials}}}}$
Here, $f_{\text{comparison},i}$ and $f_{\text{QUASAR},i}$ are the final error values achieved by the competitor algorithm (e.g., DE or L-SHADE) and QUASAR, respectively, in the $i$-th trial. GMERF>1{GMERF}>1 indicates that QUASAR performed better than the comparison optimizer.
For the final aggregated results, the Overall $GMERF$ is calculated as the geometric mean of the GMERF values obtained for each function/dimension combination.
$\text{GMERF}_{\text{Overall}}=\left(\prod_{j=1}^{N_{\text{combinations}}}\text{GMERF}_{j}\right)^{\frac{1}{N_{\text{combinations}}}}$
Computational Efficiency:

### v3

@@ Abstract.
High-dimensional numerical optimization presents a persistent challenge in computational science. This paper introduces Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), an evolutionary algorithm to accelerate convergence in complex, non-differentiable problems afflicted by the curse of dimensionality.
QUASAR expands upon the core principles of Differential Evolution (DE), introducing quasi-adaptive mechanisms to dynamically balance exploration and exploitation in its search. Inspired by the behavior of quantum particles, the algorithm utilizes three highly stochastic mechanisms that augment standard DE: 1) probabilistic mutation strategies and scaling factors; 2) rank-based crossover rates; 3) asymptotically decaying covariance reinitializations.
Evaluated on the notoriously difficult CEC2017 benchmark suite of 29 test functions, QUASAR achieved the lowest overall rank sum (367) using the Friedman test, outperforming DE (735) and L-SHADE (452). Geometric mean comparisons show average final solution quality improvements of $3.85\times$ and $2.07\times$ compared to DE and L-SHADE, respectively ($p\ll 0.001$), with average optimization speed averaging $1.40\times$ and $5.16\times$ faster.
QUASAR’s performance establishes it as an effective, efficient, and user-friendly evolutionary algorithm for complex high-dimensional problems.
@@ 1. Introduction
High-dimensional numerical optimization remains a persistent challenge across many fields, including machine learning, science, engineering, and finance (Zhou et al., 2024). This challenge stems from the curse of dimensionality; the exponential growth of the search space renders gradient-based and parametric search methods ineffective (Shan and Wang, 2008). Consequently, optimization metaheuristics, particularly evolutionary algorithms, have become essential for locating optimal solutions in complex function landscapes.
Among these methods, Differential Evolution (DE) has established itself as a robust and highly effective optimizer due to its simplicity, efficiency, and effective mutation strategies (Storn, 1996). However, standard DE variants can struggle to maintain an effective balance between global exploration and local exploitation, particularly in high-dimensional search spaces. This often leads to stagnation or premature convergence that requires frequent restarts or extensive hyperparameter tuning, a significant practical drawback for real-world applications.
To address these limitations, this paper proposes Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR), an evolutionary algorithm designed to dynamically balance the exploration-exploitation trade-off and achieve superior performance in high-dimensional spaces. This method is conceptually inspired by the search for stability in a stellar core, where quantum particles probabilistically explore configurations searching for the most stable state.
@@
The efficacy of QUASAR is demonstrated through rigorous testing against standard DE and its state-of-the-art variant L-SHADE across the complete CEC2017 benchmark suite. The results show large and statistically significant performance gains; these improvements are achieved using default parameters, highlighting QUASAR’s ease of use.
The remainder of this paper is organized as follows: Section 2 details the QUASAR algorithm, including its mutation, crossover, selection, and reinitialization mechanisms. Section 3 defines the experimental design, and Section 4 analyzes the experimental results compared to DE and L-SHADE. Section 5 describes the ease of use and implementation of the algorithm. Section 6 concludes the paper, with Section 7 proposing future work.
@@ 2. QUASAR Algorithm
Quasi-Adaptive Search with Asymptotic Reinitialization (QUASAR) operates using a highly stochastic, multi-strategy approach. The mechanisms defined below synergize to create a dynamic ’push-and-pull’ effect on the solution vectors.
@@ 2.1. Initial Population
The initial population is generated as a Sobol sequence by default, due to its low discrepancy and superior uniformity in high-dimensional spaces compared to other quasi-Monte Carlo sampling methods (Bonneel et al., 2025). The default population size, $N$, is set to $10\times D$, where $D$ is the number of dimensions.
Additional initialization methods include uniform Latin Hypercube and random distributions, as well as non-uniform Hyperellipsoid Density sequences (Soltes, 2025).
@@ 2.2. Mutation
Solutions are recombined with one another, analogous to particles sharing information via quantum entanglement. This ’spooky action at a distance’ (A. Einstein, 1947) serves as the conceptual basis for the mutation mechanisms below.
QUASAR uses three distinct mutation strategies, each being a modified Differential Evolution variant. These strategies are selected dynamically based on the specified entanglement probability, entangle rate. This value defaults to $0.33$, resulting in the three strategies being applied equally across the population.
2.2.1. Local Mutation (Exploitation):
The local mutation strategy, ’Spooky-Best’, is chosen with probability entangle rate. This strategy is similar to the DE/best/1 format, exploiting the immediate search space around the current best solution $\mathbf{X}_{\text{best},g}$.
@@ 2.2.1. Local Mutation (Exploitation):
2.2.2. Global Mutations (Exploration):
Two global mutation strategies are applied to the remaining $(1-\textit{entangle rate})$ fraction of the population. Promoting global exploration, these strategies, ’Spooky-Current’ and ’Spooky-Random’, each have a 50% probability of being selected:
@@ 2.2.2. Global Mutations (Exploration):
  • $\mathbf{v}_{i,g}$: Mutant vector for the $i$-th individual at generation $g$.
  • $\mathbf{X}_{i,g}$: Current solution vector.
  • $\mathbf{X}_{\text{best},g}$: Best known solution in the population.
  • $\mathbf{X}_{\text{rand},g}$: Randomly chosen solution vector.
  • $F_{\text{local}}$ and $F_{\text{global}}$: Corresponding mutation factors.
@@ 2.3. Crossover
QUASAR uses binomial crossover with a dynamic crossover rate ($CR_{i}$) proportional to the solution’s fitness rank. The best solutions are rapidly recombined to maximize both exploration and exploitation via synergy with global and local mutations.
The raw crossover rate ($CR_{\text{raw}},i$) for the $i$-th solution is calculated based on its rank ($\textit{rank}_{i}\in[0,N-1]$), where $\textit{rank}=0$ corresponds to the best solution (lowest objective function value). The final dynamic crossover rate $CR_{i}$ is capped at a minimum of $0.33$. This mechanism ensures that the best solutions have a higher chance of inheriting components from the new mutant vector $\mathbf{v}_{i}$.
(1) $CR_{\text{raw},i}=\frac{N-1-\text{rank}_{i}}{N-1}$
(2) $CR_{i}=\max(CR_{\text{raw},i},0.33)$
@@ 2.4. Selection
A greedy elitism strategy is used to determine the next generation’s population, where the target vector for generation $g+1$ is the better of the two solutions between the current vector $\mathbf{X}_{i,g}$ and the trial vector $\mathbf{u}_{i,g}$. If a solution undergoes reinitialization, it bypasses the selection step.
@@ 2.5. Asymptotic Reinitialization
The covariance-guided vector reinitialization applies to a fixed 33% of the worst-performing solutions. In early generations, these solutions have a high probability of ’tunneling’ to more promising regions. This mechanism complements the rank-based crossover; while the best solutions recombine frequently, lower quality solutions explore locally until stochastically tunneling to a better location.
@@
(5) $P_{\text{reinit}}(g)=\exp\left(\frac{\ln(P_{\text{final}})}{g_{\text{final}}\cdot g_{\text{max}}}\cdot g\right)$
where $g_{\text{max}}$ is the maximum generation count, $P_{\text{final}}=0.33$ is the target final probability, and $g_{\text{final}}=0.33$ defines the fraction of generations at which $P_{reinit}$ reaches $P_{final}$. This ensures the probability drops to $P_{\text{final}}$ at $g=g_{\text{max}}\cdot g_{\text{final}}$.
During a reinitialization event, new solutions are sampled from a multivariate distribution shaped by top performing individuals. The top $M$ solutions (where $M=25\%\text{of the population}$) are used to calculate the mean vector $\boldsymbol{\mu}$ and the covariance matrix $\boldsymbol{\Sigma}$:
@@
(7) $\boldsymbol{\Sigma}=\frac{1}{M-1}\sum_{i=1}^{M}(\mathbf{X}_{i}-\boldsymbol{\mu})(\mathbf{X}_{i}-\boldsymbol{\mu})^{\top}+\epsilon\mathbf{I}$
where $\epsilon$ is a small scalar ($10^{-12}$) and $\mathbf{I}$ is the identity matrix, added to prevent $\boldsymbol{\Sigma}$ from being singular. Similarly, the asymptotic reinitialization is bypassed for cases where $N<D$. The effects of skipping the reinitialization can be seen in the $N=100$ sample-variant trials (Fig. 4).
@@
The noise vector $\boldsymbol{\delta}$ is sampled from a multivariate normal distribution, with respect to the search bounds:
(9) $\boldsymbol{\delta}\sim\mathcal{N}\left(0,\left(\frac{\mathbf{b}_{\text{high},n}-\mathbf{b}_{\text{low},n}}{20}\right)^{2}\mathbf{I}\right)$
@@
(10) $\mathbf{X}_{\text{new},n}=\text{clip}(\mathbf{y}_{n},\mathbf{B}_{n})$
@@
Due to the covariance calculations only using a fraction of the population, along with decaying probability of occurrence, the additional overhead becomes negligible compared to the objective function evaluations. For extreme cases, the reinitialization can be disabled, which is the default behavior for $N<D$ problems.
@@ 3. Experimental Design
3.1. Optimizers & Test Functions
@@
This choice is justified by the fact that the QUASAR algorithm is built on the DE framework, as well as DE’s standing as a robust and effective optimizer (Price, 1996). L-SHADE was specifically chosen due to its success in CEC optimization competitions (Mo et al., 2019).
@@ 3.2. Experimental Setup
To ensure a fair comparison, each optimizer was run using their default hyperparameters. The initial population generation methods were also left to defaults; the tested sample sizes not being powers of 2 diminishes the uniformity of QUASAR’s initial Sobol sequence (Joe and Kuo, 2008), providing a theoretical initial handicap.
Since SciPy’s DE implementation internally scales the input population size by dimensionality ($N\times D$ when all dimensions have equal bounds), a correction factor of $1/D$ is applied to its initial sample size.
The computational budget for metaheuristic comparison is often defined by the maximum number of objective function evaluations ($FES_{max}$) (Kazikova et al., 2021). Due to L-SHADE’s linear population reduction, the $FES_{max}$ becomes a difficult comparison metric to compare against. As such, the primary comparisons are performed on final solution fitness errors and overall optimization times, given the same number of generations, using modern implementations via SciPy’s DE and MealPy’s L-SHADE.
By default, the L-SHADE implementation features tolerance parameters for early convergence. Additionally, the QUASAR and DE implementations include a final polishing step, using Powell and L-BFGS-B minimization, respectively. These were all disabled for the trials to ensure the optimizers fully utilize their computational budget, terminating only when reaching the maximum number of generations $g_{max}$.
@@ 3.3. Experimental Trials:
The results are compared based on the overall aggregated rank sums, geometric mean error factors ($R_{\text{error}}$), and optimization run time ratios $R_{\text{time}}$.
3.3.1. Dimension-Variant Trials:
The dimension-variant trials assessed QUASAR’s scalability by scaling the problem dimension $D$ with a fixed population size ($N=1000$). $30$ independent trials were performed for each test function/dimension combination to ensure statistical significance.
The dimensions were varied between $[10,30,50,100]$, resulting in a total of 10,400 full optimization runs; 3,480 for each algorithm. The computational budget was kept constant, with maximum number of iterations $g_{max}=100$ and fixed sample size $N=1000$.
3.3.2. Sample-Variant Trials:
The sample-variant trials assessed the impact of population size on performance. The computational budget is varied between population sizes of $[100,250,500,1000,5000]$. $30$ independent trials were performed for each test function / sample size combination, using maximum number of iterations $g_{max}=100$ and fixed dimensionality $D=100$.
The $N=5000$ trials, however, were only run for 10 trials to ensure statistical significance while mitigating the high computational overhead, resulting in a total of 11,310 full optimization runs; 3770 for each optimization algorithm.
Table 1. Experimental Parameters for Sample and Dimension Variations
Sample-variant Dimension-variant
D 100 [10, 30, 50, 100]
N [100, 250, 500, 1000, 5000] 1000
$\mathbf{g_{\max}}$ 100 100
@@ 3.4.1. Overall Rank:
For each unique combination, the median final error value $f$ across all independent trials was calculated for each optimizer. The optimizers were ranked based on these median errors, with the best performance (lowest error) receiving a rank of 1.
@@
3.4.2. Solution Quality Improvements:
The primary comparison metric used to quantify the magnitude of performance improvement is the geometric mean error reduction factor $R_{\text{error}}$. This provides an unbiased measure of relative final objective penalty $f$ for the comparison optimizer to QUASAR across all $T$ independent trials for each scenario:
$R_{\text{error}}=\sqrt[T]{\prod_{i=1}^{T}\frac{f_{\text{comp},i}}{f_{\text{QUASAR},i}}}$
Here, $f_{\text{comp},i}$ and $f_{\text{QUASAR},i}$ are the final error values achieved by the competitor algorithm (e.g., DE or L-SHADE) and QUASAR, respectively, in the $i$-th trial.
For the overall results, the average $\bar{R}_{\text{error}}$ across all data is calculated as the geometric mean of the $R_{\text{error}}$ obtained for each of $k$ function/dimension or function/sample size combinations.
${\bar{R}_{\text{error}}}=\sqrt[k]{\prod_{j=1}^{k}{R_{\text{error},j}}}$
3.4.3. Computational Efficiency: